OpenStudy (anonymous):

What is the derivative of (2x)^(2x)?

zepdrix (zepdrix):

Ooo fun problem :o

zepdrix (zepdrix):

We have a variable in the exponent, so we can't apply power rule. We have a variable base, so we can't apply exponential rule. What shall we do!! :)

OpenStudy (anonymous):

Something to do with log..?

zepdrix (zepdrix):

Yessss, something with log:)

OpenStudy (anonymous):

Uhhh...... I would say take the log of both sides but there's nothing on the other side :/

zepdrix (zepdrix):

We can `make up` another side if you want. Say y=(2x)^(2x). But there is another trick we can use. It's really important to get comfortable with it if you're going to go on to higher math classes.

OpenStudy (anonymous):

What's the trick?

zepdrix (zepdrix):

So there is this idea that like... \[\Large\rm e^{\ln(x)}=x\]If you take the composition of inverse functions, they "undo" one another. So as I've posted above, this look familiar?

zepdrix (zepdrix):

We want to use this idea in reverse.\[\Large\rm \color{orangered}{\left[(2x)^{2x}\right]}=e^{\ln\color{orangered}{\left[(2x)^{2x}\right]}}\]Or does this approach look too scary? :)

OpenStudy (anonymous):

Don't the e and the ln cancel out because ln is the log of e?

zepdrix (zepdrix):

Yes exactly, log base e, exponential function base e. Cancel out isn't quite the right word. But if you want to think of it in those terms, that's fine. :)

OpenStudy (anonymous):

Okay, so then what do I do?

zepdrix (zepdrix):

Don't they cancel out?* They could, yes. But we don't want to cancel them out. We want to use that jumbled mess because it's easier to differentiate.

OpenStudy (jdoe0001):

hmm \(\bf (2x)^{2x}\implies (2x)^x\cdot (2x)^x\)

zepdrix (zepdrix):

\[\Large\rm \frac{d}{dx}e^{stuff}\]You remember how to differentiate exponential base e?

OpenStudy (anonymous):

Uhh isn't it e^u(u')?

zepdrix (zepdrix):

\[\Large\rm \frac{d}{dx}e^{stuff}=e^{stuff}(stuff)'\]Yes good.

zepdrix (zepdrix):

\[\Large\rm \frac{d}{dx}e^{\ln\left[(2x)^{2x}\right]}=e^{\ln\left[(2x)^{2x}\right]}\frac{d}{dx}\ln\left[(2x)^{2x}\right]\]So we get something like that, yah? Same exponential back. Just need to multiply by derivative of the log.

zepdrix (zepdrix):

The whole reason for doing this ... is that logs allows us to get variables out of the exponent position. Remember your log rule? :o

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

*still confused*

zepdrix (zepdrix):

Too much? :c Should we go back to the normal method you're used to? Get lost in this method?

OpenStudy (anonymous):

Kind of D:

zepdrix (zepdrix):

\[\Large\rm y=(2x)^{2x}\]Taking the log of each side,\[\Large\rm \ln y=\ln (2x)^{2x}\]Ok applying our log rule,\[\Large\rm \ln y= 2x \ln(2x)\]From here we differentiate each side with respect to x. Looks like we have to product rule on the right side. Understand how to differentiate the left side?

OpenStudy (anonymous):

(1/y) = (2x)(2/2x)+(2)(ln(2x))?

zepdrix (zepdrix):

One thing to fix, \[\Large\rm \frac{1}{y}\color{red}{y'}=2x\frac{2}{2x}+2\ln2x\]We're differentiating some variable with respect to x, so a y' should pop out, yes?

OpenStudy (anonymous):

Yes, and then solve for y'?

zepdrix (zepdrix):

yah :)

zepdrix (zepdrix):

recall that \(\Large\rm y=(2x)^{2x}\) So we'll need to plug that in for y at some point.

OpenStudy (anonymous):

Got it!! Thanks so much :)